OpenStudy (anonymous):

A vending machine's coin box contains nickels, dimes, and quarters. The total number of coins in the box is 336. The number of dimes is three time the number of nickels and quarters together. If the box contains 33 dollars and 20 cents, find the number of nickels, dimes, and quarters that it contains.

OpenStudy (mertsj):

Let n = the number of nickels Let d = the number of dimes Let q = the number of quarters What is n+d+q ?

OpenStudy (anonymous):

n+d+q = 336 3(n+q) = d

OpenStudy (mertsj):

n nickels are worth 5n cents or .05n dollars d dimes are worth 10d cents or .10d dollars q quarters are worth 25q cents or .25 dollars

OpenStudy (mertsj):

Can you write the third equation now?

OpenStudy (anonymous):

.05n + .10d + .25q = 33.20 This is where i get stuck

OpenStudy (mertsj):

I would make my life easier and multiply by 100 to get 5n+10d+25q=3320

OpenStudy (mertsj):

Then I would divide by 5 to get n + 2d+5q=664

OpenStudy (mertsj):

Does that help?

OpenStudy (anonymous):

not really i am using substitution method be not getting any where (n + 2d+5q=664 +( -1)n+d+q = 336) = d+4q=328 ???

OpenStudy (mertsj):